### v1

# Can Global XAI Methods Reveal Injected Bias in LLMs? SHAP vs Rule Extraction vs RuleSHAP

Abstract
Generative AI systems can help spread information but also misinformation and biases, potentially undermining the UN Sustainable Development Goals (SDGs). Explainable AI (XAI) aims to reveal the inner workings of AI systems and expose misbehaviours or biases. However, current XAI tools, built for simpler models, struggle to handle the non-numerical nature of large language models (LLMs). This paper examines the effectiveness of global XAI methods, such as rule-extraction algorithms and SHAP, in detecting bias in LLMs. To do so, we first show a text-to-ordinal mapping strategy to convert non-numerical inputs/outputs into numerical features, enabling these tools to identify (some) misinformation-related biases in LLM-generated content. Then, we inject non-linear biases of varying complexity (univariate, conjunctive, and non-convex) into widespread LLMs like ChatGPT and Llama via system instructions, using global XAI methods to detect them. This way, we found that RuleFit struggles with conjunctive and non-convex biases, while SHAP can approximate conjunctive biases but cannot express them as actionable rules. Hence, we introduce RuleSHAP, a global rule extraction algorithm combining SHAP and RuleFit to detect more non-univariate biases, improving injected bias detection over RuleFit by +94% (MRR@1) on average.
1 Introduction
The United Nations’ Sustainable Development Goals (SDGs) address global challenges such as poverty (SDG 1), education (SDG 4), clean water (SDG 6), inequality (SDG 10), and climate change (SDG 13), each with specific targets for a sustainable future [1]. Meanwhile, AI has the potential to both accelerate and decelerate progress toward these goals [2]. For instance, [3] found limitations in the depth and reliability of generative AI for climate communication (SDG 13).
While large language models (LLMs) are widely used for knowledge discovery and education, their potential to spread biased or inaccurate information poses risks [4]. For instance, an LLM might dismiss a promising technology, such as the use of micro-algae for CO2 absorption, as ineffective, deterring users or investors from pursuing such solutions solely based on an LLM’s misrepresentation. Hence, to ensure AI contributes positively to the SDGs, it is critical to improve the quality of generated information in widespread LLM-powered applications. This will help align, among other things, AI technologies with the global sustainability agenda.
The complexity of LLM training pipelines [5, 6], coupled with the lack of publicly available training data [7, 8], makes it hard to analyse these models and detect biases arising from skewed or polluted datasets. Moreover, biases can be introduced (whether unintentionally or maliciously) through system instructions or other simple methods, especially by third-party users deploying ChatGPT or similar services in their applications. Consequently, identifying bias in LLMs requires a generic, black-box-compatible approach, motivating the use of explainable AI (XAI).
However, existing XAI tools, originally designed for simpler AI systems, struggle to scale to state-of-the-art LLMs [9], limiting our ability to detect biases in contexts like the UN SDGs. For example, SHAP [10] uses Shapley values to identify influential input features. While robust at the global scale, it relies on numerical inputs and ordinal outputs, unlike the vocabulary-based input/output space of LLMs. In contrast, RuleFit [11] uses gradient boosting [12] to extract rule-based relationships and LASSO [13] to prune them. Although RuleFit supports both categorical and ordinal spaces, it lacks SHAP’s theoretical guarantees and still struggles with the input/output space of LLMs.
To investigate global bias in widespread LLMs (i.e., ChatGPT, Llama) and detect biases, we first introduce text-to-ordinal mapping, a strategy to abstract the input and output spaces of LLMs into numerical, ordinal features, enabling the application of methods like SHAP and RuleFit. Inspired by the recent LLM-as-a-judge paradigm [14], our ordinal features are specifically crafted to reflect LLMs’ implicit judgments (i.e., opinions) on text properties linked to cognitive biases. These include phenomena such as framing effects, information overload, and oversimplification, key factors in the dissemination of misinformation on social networks [15, 16].
Hence, we present an automated assessment of global XAI algorithms, evaluating their ability to detect that kind of opinion-based biases in LLMs. Our assessment is based on the manual injection of biases via system instructions (preliminary prompts that shape how a LLM responds). We introduce three levels of bias complexity (univariate threshold, conjunctive multi-threshold, and non-convex/disconnected) to test the algorithms’ performance under varied conditions. We then evaluate how many of these rules are detected by state-of-the-art XAI methods with public implementation.
While SHAP often identifies the injected biases, it cannot express them as actionable, symbolic rules. Instead, methods like RuleFit produce too many rules and often fail to pinpoint non-univariate biases precisely. To address these shortcomings, we propose RuleSHAP, a novel algorithm that integrates SHAP into RuleFit by refining rule extraction and LASSO regression. Our results show that RuleSHAP outperforms all baselines in detecting non-linear biases injected into state-of-the-art LLMs. The replication package is available at: https://github.com/Francesco-Sovrano/RuleSHAP.
2 Related Work
Few model-agnostic XAI methods can extract global, non-linear rules from ordinal inputs/outputs (cf. Section 5), and all of them rely on tree surrogates. To our knowledge, RuleSHAP is the first to integrate global SHAP values to guide model-agnostic rule generation in tree surrogates. Indeed, while [17, 18, 19] produce global explanations, they do not support symbolic rule derivation.
Recent studies explore biases in LLMs using mixed methods [3, 20, 21] or focusing on demographic fairness [22]. For instance, [23] assess ChatGPT’s SDG literacy, and [24] examine human–LLM alignment challenges. Both emphasize attitudinal alignment, contrasting with our focus on detecting and ranking injected biases. [3], instead, highlight generative AI’s limitations in climate communication (SDG 13) only. In contrast, our framework automates data collection and bias analysis, supporting extensibility to new LLMs and abstraction schemes.
The most closely related work is probably [25], which also uses LLM-as-a-judge for bias detection but focuses on gender bias, validating that these metrics align with human judgments on response bias but without resorting to explicit rule extraction. Other works address cognitive or domain-specific biases without adopting global rule-extraction for LLMs. [26] examine anchoring and framing biases, proposing self-help interventions, while [27] study LLMs as text-quality judges. In the medical domain, [28] introduce BiasMedQA to quantify clinical biases.
3 Text-to-Ordinal Mapping for Bias Detection
Preliminaries on global XAI. Two model-agnostic XAI approaches motivate our work: (i) perturbation-basedand (ii) rule-extractionmethods. Perturbation-based methods, e.g., LIME [29] and SHAP [10], analyse model responses to perturbed inputs $x^{\prime}$ generated by modifying subsets of features in $x\in\mathbb{R}^{d}$. SHAP estimates the feature importance $\rho_{i}$ to the output so as to satisfy properties like additivity: $f(x)=\sum_{i=1}^{d}\rho_{i}$. Rule-extraction methods, such as RuleFit [11], derive instead interpretable rules $R=\{r_{1},...,r_{k}\}$ from data points $X$ using surrogate models $f_{s}$ (e.g., decision trees), approximating the original model $f$ via $f_{s}(X)\approx f(X)$. Both approaches generally assume: (i) numerical or discrete input features suitable for perturbation; (ii) ordinal/categorical outputs.
Focus on bias detection within explanations. We seek to reveal biases that may affect how widespread general-purpose LLMs explain critical SDG themes. By focusing on topics where accurate, unbiased information is crucial, we can detect high-level biases that play a central role in the dissemination of misinformation [15, 16], such as framing effects (emphasizing certain attributes of an issue while downplaying others), information overload (producing overly long or complex outputs), and oversimplification (providing superficial explanations for complex matters).
Formally, let $\mathcal{T}=\{\tau_{1},\tau_{2},...,\tau_{m}\}$ be a set of topics (i.e., textual phrases), each tied to an SDG-related challenge or solution (e.g., “micro-algae for CO2 absorption”). Let $\mathcal{X}={\mathbf{x}_{1},\mathbf{x}_{2},...,\mathbf{x}_{N}}\subset\Sigma^{*}$ denote a set of prompt texts, each being a request for an explanation of a topic $\tau_{i}\in\mathcal{T}$ generated by a function $q\colon\mathcal{T}\to\mathcal{X}$ such that $\mathbf{x}_{i}=q(\tau_{i})$, where $q(\tau_{i})=$Explain $\tau_{i}$}” is the simplest choice. Then, $f:\mathcal{X}\to\Sigma^{*}$ models an LLM producing a textual explanation $\mathbf{y}_{i}=f(q(\tau_{i}))$ for each topic $\tau_{i}$.
Abstraction setting. While a naive approach could rely on a vocabulary-based encoding that treats text as categorical tokens, it would fail to capture corpus-level biases. To address this limitation in the context of bias detection, we propose extracting domain-relevant properties from both the topics and their corresponding explanations. We define two mappings, $\phi_{x}\colon\mathcal{T}\to\mathbb{R}^{d}$ and $\phi_{y}\colon\Sigma^{*}\to\mathbb{R}^{m}$, that transform textual data into numerical (ordinal) representations: $\mathbf{u}_{i}=\phi_{x}(\tau_{i})\in\mathbb{R}^{d}$ and $\mathbf{v}_{i}=\phi_{y}(\mathbf{y}_{i})\in\mathbb{R}^{m}$. Once these mappings are applied, the textual prompts and responses ($\mathbf{x}_{i}$, $\mathbf{y}_{i}$) are no longer used in the subsequent XAI pipeline. Under this framework, perturbation-based or rule-extraction XAI methods operate directly on the numeric vectors $\{\mathbf{u}_{i},\mathbf{v}_{i}\}$, thereby facilitating bias detection and interpretability in a LLM.
Main assumptions and important workaround. The mapping process assumes that certain conceptual properties are learned in an AI’s latent space to model text generation. For instance, topics like war, disease, and famine often carry a negative bias in human writing, and we expect an LLM trained on such texts to share this bias. So, if bias relates to these properties, text-to-ordinal mapping can reveal the nature of this relationship. Moreover, the framework also assumes that topics with identical $\mathbf{u}$-vectors ($\mathbf{u}_{i}=\mathbf{u}_{j}$) yield similar output vectors $\mathbf{v}_{i}$ and $\mathbf{v}_{j}$, an assumption that may fail due to LLM non-determinism. To address this, a sufficiently large $\mathcal{T}$ is needed to ensure statistical power and detectable correlations between $\mathbf{u}_{i}$ and $\mathbf{v}_{i}$. These correlations can be estimated via Pearson or Spearman methods for linear or monotonic biases, or via distance correlation [30] for non-linear biases (as in our case). Thus, any global XAI insight from text-to-ordinal mappings should be supported by correlation evidence certifying abstraction validity.
Input abstractions. Our analysis focuses on topics $\tau_{i}\in\mathcal{T}$ relevant to the UN SDGs. Recall that each topic $\tau_{i}\in\mathcal{T}$ is mapped to an ordinal feature vector $\mathbf{u}_{i}=\phi_{x}(\tau_{i})\in\mathbb{R}^{d}$. In a standard LLM scenario, we do not have direct access to the training corpus. Hence, we propose using the LLM itself as $\phi_{x}(\cdot)$: as a judge providing informed opinions about a topic. In this way, an XAI algorithm can summarise how LLM’s explicit opinions on topic properties reflect underlying (opinion-based) biases. To assess whether different types of topics elicit skewed explanations, we define a set of 11 input properties mapped by $\phi_{x}(\cdot)$. These capture how the LLM perceives factors such as a topic’s prevalence in public discourse (commonality), emotional tone (positivity, negativity, neutrality), and breadth across fields (interdisciplinary nature). Additional properties include conceptual complexity, controversy, and regional and temporal variation. For a detailed breakdown of all properties, the specific prompts used to elicit LLM opinions, and other hyperparameters, see Appendix A. Specifically, LLMs are asked to rate, on a scale from 1 (low) to 5 (high), how a topic is typically portrayed on the web, e.g., whether it is common, positive, controversial, etc. Thus, for each topic $\tau_{i}$, we obtain a vector $\mathbf{u}_{i}=\phi_{x}(\tau_{i})\in\mathbb{R}^{d}$ ($d=11$), where each component is an ordinal score on a 1–5 Likert scale.
Output abstractions. Given a set of LLM-generated textual responses $\mathbf{y}_{i}=f(q(\tau_{i}))$, we define 7 output features to quantify relevant properties of each explanation $\mathbf{y}_{i}$. All these features are mapped to numerical scores by $\phi_{y}(\cdot)$. Thus, for each output $\mathbf{y}_{i}$, we construct a vector $\mathbf{v}_{i}=\phi_{y}(\mathbf{y}_{i})\in\mathbb{R}^{m}$ ($m=7$) capturing these properties. The first four features are proxy metrics commonly associated with the biases under study: (i) Explanation length (characters count) reflects information overload (if high) or oversimplification (if low) [31, 32]. (ii) Subjectivity and (iii) sentiment scores capture subjectivity and polarity, highlighting framing effects [33]. (iv) Readability indices, such as the Gunning fog index [34], may suggest instead information overload or oversimplification. The corresponding scoring functions $\phi_{y}(\cdot)$ are: a well-known readability index (Gunning fog index) yielding continuous scores [34]; character count as the measure of explanation length; publicly available sentiment/subjectivity estimators to generate sentiment/subjectivity scores (for more details see Appendix D). The remaining three features directly estimate (v) information overload, (vi) oversimplification, and (vii) framing effects. Inspired by [25], for these features we follow the strategy used with $\phi_{x}(\cdot)$, asking the LLM to rate (from 1 to 5) how much an explanation can produce a given bias (see Appendix B). This approach helps cross-validate whether our text-to-ordinal mapping captures bias-related signals. Indeed, if there is a bias, the direct estimates are expected to correlate with the four proxy metrics.
4 RuleSHAP
We hereby outline SHAP and RuleFit. The implementation details of RuleSHAP will also follow.
SHAP [10] is a model-agnostic XAI framework based on cooperative game theory. Given a model $f$ and an instance $u$, SHAP assigns each feature $u_{i}$ a Shapley value $\rho_{i}$, representing its average marginal contribution to the prediction $f(u)$ across all feature coalitions. This is estimated by replacing features with predefined background values (which are not necessarily 00) to measure the impact of their removal. For global attributions, SHAP averages local attributions over multiple instances [35], providing an overall measure of feature importance. These values are often visualized in summary plots like bar charts or beeswarm plots. Since computing exact Shapley values is costly and assumes feature independence, practical approximations like KernelSHAP, which employs weighted linear regression to approximate contributions, or permutation-based methods that limit feature permutations to second-order interactions, i.e., the permutation SHAP explainer (used in our experiments). To mitigate independence assumption errors, correlated features can be grouped before computation.
RuleFit [11] is a global XAI method that extracts the rules a model $f$ follows to produce outputs $\mathbf{v}$ from inputs $\mathbf{u}$. It does this by leveraging features within $\mathbf{u}$, where each input vector consists of multiple features $c_{1},c_{2},\dots,c_{m}$. RuleFit combines gradient boosting [12] with sparse linear modelling to build interpretable models. Gradient boosting iteratively constructs decision trees, where each split minimizes impurity or maximizes loss reduction. A decision path from root to leaf defines a binary rule, e.g., $(c_{i}\leq\theta_{1})\land(c_{j}>\theta_{2})\land\dots$, where $\theta_{1},\theta_{2},\dots$ are threshold values. After extracting rules, RuleFit forms a binary design matrix $\mathbf{X}$, with rows as instances and columns as rules. Each entry $X_{k,n}$ is 1 if the $k$-th rule is activated for the $n$-th instance, otherwise 0. $\mathbf{X}$ has dimensions $|R|\times|\mathbf{u}|$ (where $R$ is the set of extracted rules), capturing the presence of each rule across all instances. Then, given the coefficient vector $\mathbf{w}$ and an hyper-parameter $\alpha$ to control for sparsity, RuleFit applies a LASSO regression to select a sparse set of predictive rules by solving: $\min_{\mathbf{w}}\ \frac{1}{2}\|\mathbf{v}-\mathbf{X}\mathbf{w}\|_{2}^{2}+\alpha\|\mathbf{w}\|_{1}.$
Technical Challenges and Proposed Solution. SHAP provides rigorous game-theoretic feature attributions with additivity and consistency but lacks global symbolic representations, limiting its utility when human-readable rules are needed, particularly for complex relationships. RuleFit generates interpretable rules with gradient-boosted trees and LASSO but lacks SHAP’s theoretical guarantees, so its rules may not reflect a truly global measure of feature importance. To overcome the limitations of SHAP and RuleFit, we introduce RuleSHAP, a novel algorithm integrating SHAP into RuleFit. This integration leverages SHAP’s robust feature attribution, preserving RuleFit’s interpretability and rule-based representation. To achieve this, we follow these steps:
Step 1: Global Shapley Values Aggregation. We first compute SHAP values for all available data points using one of the SHAP explainers available in the official repository. Specifically, for each feature $c_{i}$ and instance $\mathbf{u}_{k}$, we obtain the Shapley value $\rho_{(i,k)}$, using the minimum feature values from $\mathcal{T}$ as the background since all features are, by construction, ordinal. Because we work with abstracted features, SHAP cannot directly estimate the average marginal contribution $\rho_{(i,k)}$, by simple feature “removal”, sending a new $\widehat{\mathbf{u}}_{k}$-vector to the LLM for evaluation.
Instead, following the workaround from Section 3, our strategy is to guarantee sufficient redundancy in $\mathcal{T}$ so that we can find multiple points $j$ in $\mathcal{T}$ for which $\|\widehat{\mathbf{u}}_{k}-\mathbf{u}_{j}\|_{2}$ is minimal ($\approx 0$) while matching the background values considered by SHAP perturbations. SHAP can then randomly select one of these points for contribution estimation, mimicking the non-determinism of LLMs and enabling a statistical assessment of how input abstractions correlate with bias-related properties. Specifically, we aggregate the $\rho_{(i,k)}$ values across all instances by calculating the mean of the absolute Shapley values, adding the standard deviation to capture the upper bound of each feature’s importance:
$\rho_{i}^{\text{agg}}=\frac{1}{N}\sum_{k=1}^{N}|\rho_{(i,k)}|+\text{std}\left(\{|\rho_{(i,k)}|\}_{k=1}^{N}\right)$
where $N$ is the total number of data points $\mathbf{u}$. The aggregated Shapley values $\rho_{i}^{\text{agg}}$ are then normalized to the interval $(0,1]$ by: $\bar{\rho_{i}}=\rho_{i}^{\text{agg}}/\sum_{j=1}^{d}\rho_{j}^{\text{agg}}$, where $d$ is the number of input features.
Step 2: XGBoost with SHAP Weighting. In this step, we enhance the RuleFit model by replacing its traditional gradient boosting component with XGBoost [36, eXtreme Gradient Boosting], an optimized implementation of gradient boosting that introduces regularization, parallel processing, and efficient handling of missing data. In particular, XGBoost allows for feature weighting during rule selection, where to each feature $c_{i}$ is possible to assign a weight, which in our case is derived from the aggregated Shapley values $\bar{\rho_{i}}$ computed in Step 1.
To integrate SHAP-driven rule selection into XGBoost, we have to adjust specific hyperparameters to leverage feature weights effectively. First, we configure colsample_bylevel to $\frac{1}{M}$, which controls the fraction of features sampled at each tree level. By setting this parameter accordingly, the model is encouraged to sample approximately one feature per level. Specifically, this sampling process is guided by the computed feature weights $\{\bar{\rho_{i}}\}$, meaning that features with higher Shapley values have a greater likelihood of being selected at each level. This prioritization ensures that globally important features play a more significant role in model construction. Additionally, we employ the exact setting for the tree_method hyperparameter, which directs XGBoost to use a precise greedy algorithm that thoroughly evaluates all possible split points across selected features. This exhaustive approach guarantees the identification of optimal splits at each node, minimizing the risk of suboptimal partitioning that could otherwise lead to underfitting due to the low colsample_bylevel value.
Step 3: SHAP-Aware LASSO Regression. Finally, we modify the LASSO regression component to prioritize rules that involve features with higher global aggregated Shapley values. For each extracted rule $r_{j}$, identify the set of features $\mathcal{F}(r_{j})$ that appear in its conditions. Compute the average aggregated Shapley value for these features: $\rho_{r_{j}}=\frac{1}{|\mathcal{F}(r_{j})|}\sum_{x_{i}\in\mathcal{F}(r_{j})}\bar{\rho_{i}}$. These rule-specific weights $\rho_{r_{j}}$ reflect the overall importance of the features involved in each rule, prioritizing rules with fewer and more important features. We then adjust the LASSO regularization term to incorporate these weights, modifying the optimization problem as follows: $\min_{\mathbf{w}}\ \frac{1}{2}\|\mathbf{v}-\mathbf{X}\mathbf{w}\|_{2}^{2}+\alpha\sum_{j=1}^{|R|}\frac{|\mathbf{w}_{j}|}{\rho_{r_{j}}}.$
By scaling the regularization term inversely with $\rho_{r_{j}}$, rules associated with more important features are less penalized, thereby being more likely to be retained in the final model. This approach ensures that the selected rules not only capture significant feature interactions but also align with the global importance measures provided by SHAP.
Step 4: Rule Importance Computation. Additionally, differently from RuleFit, we compute the importance of a rule $r_{j}$ using only the absolute value of its coefficient learned via LASSO regression, i.e., $I(r_{j})=|w_{j}|$, where $w_{j}$ is the coefficient assigned to rule $r_{j}$ by the LASSO model. In contrast, RuleFit defines rule importance as $I_{\text{RuleFit}}(r_{j})=|w_{j}|\cdot\sqrt{\text{support}(r_{j})\cdot(1-\text{support}(r_{j}))}$, where $\text{support}(r_{j})$ represents the proportion of training samples that satisfy rule $r_{j}$. We avoid this support-based weighting because it inherently penalizes rules with either very high or low support. This can be problematic, as some injected biases may only affect a (small) subset of instances.
5 Experimental Setup
We designed an experiment to evaluate the effectiveness of global XAI methods in detecting SDG-related biases within prominent general-purpose LLMs: (i) GPT-4o, (ii) GPT-4o mini, (iii) GPT-3.5 Turbo, (iv) Llama-3.1-70B, (v) Llama-3.1. The process (see Figure 1) begins with selecting a diverse set of topics $\mathcal{T}$ related to the UN SDGs. Then, to establish a ground truth for bias detection, we introduce biases $b_{j}$ into an LLM $f$ through system instructions, creating modified models $f_{b_{j}}$. These biases vary in complexity, incorporating: (i) univariate, (ii) conjunctivemultivariate, and (iii) non-convex/disconnected (e.g., modulus operator) relationships to simulate realistic bias scenarios. Using the instructions, we then obtain explanations $\mathbf{y}_{i}=f_{b_{j}}(\mathbf{x}_{i})$, producing a corresponding set of biased responses. These inputs and outputs are further abstracted using mapping functions $\phi_{x}(\mathbf{x}_{i})$ and $\phi_{y}(\mathbf{y}_{i})$ as defined in Section 3. To mitigate stochastic effects, we set sampling temperature and top-p to 0. Next, global XAI algorithms (e.g., RuleSHAP) are applied to each biased model $f_{b_{j}}$ to detect the injected biases. Further details on the experimental setup are provided below.
Figure 1: Experiment pipeline: Green: ground truth injection; Yellow: data auto-generation.
Topic extraction. For the topic selection phase, we focus on three domains—climate action (SDG 13), good health and well-being (SDG 3), and gender equality (SDG 5)—each corresponding to one of three distinct SDGs of the UN. These domains were chosen for their global relevance and the diverse range of topics they encompass. Within each domain, we categorize topics according to the dimensions defined in Section 3, such as whether a topic is conceptually dense, technically complicated, socially controversial, or emotionally charged, among others. Each of these dimensions is scored on a Likert scale indicating the degree to which the dimension applies to a particular topic.
Following the workaround from Section 3, we determined the required sample size for $\mathcal{T}$ through an a priori power analysis, targeting a minimum power of 0.8, a medium effect size (0.35), and $\alpha\leq 0.05$. This indicated that about 60 data points per dimension are needed. Accordingly, each LLM was tasked with generating at least n_topics = 60 unique topics per domain and score. The prompt template and hyperparameters are detailed in Appendix C. This process was repeated across the five possible scores (1–5) for each dimension to ensure comprehensive coverage. Since topics for one dimension are abstracted across others, we eventually achieve a final sample size of $60\times 11=660$ topics; well beyond the minimum required. Thus, if many similar data points (in the abstracted input space) yield outputs with the same bias-related properties, it becomes increasingly likely that their abstractions are truly correlated with these properties.
After collecting the topics, near-duplicates were removed by calculating the semantic similarity among all topics with a transformer-based model, i.e., all-MiniLM-L6-v2 [37], applying a cosine similarity threshold of 0.9 (i.e., 90% similarity). Any topic exceeding it is considered too similar and removed. This ensures that the final list (see statistics in Table 1) avoids redundancy.
Table 1: Unique topics found per LLM for each SDG domain, with total topics per LLM.
SDG ID GPT-3.5 turbo GPT-4o mini GPT-4o Llama-3.1 Llama-3.1-70B
Climate Act. 13 1,372 2,275 2,870 2,974 1,864
Well-Being 3 1,511 2,488 3,391 3,903 3,382
Gender Eq. 5 1,297 2,095 2,585 3,667 2,932
Total 4,180 6,858 8,846 10,544 8,178
Injecting bias via system instructions. Let $f$ be an LLM, and let $B=\{b_{1},b_{2},b_{3}\}$ be a set of bias-inducing system instructions of increasing non-linearity. The biased variant of $f$ under instruction $b_{j}$ is denoted $f_{b_{j}}$. It consists of a Python wrapper around $f(\mathbf{x}_{i})$ that sends instruction $b_{j}$ if the input abstraction scores produced by $\phi_{x}(\tau_{i})$ satisfy the preconditions required by $b_{j}$, where $\mathbf{x}_{i}=q(\tau_{i})$. This procedure guarantees the presence of ground truth biases we can use for assessing XAI methods.
The simplest bias condition, $b_{1}$, adjusts the length of the model’s responses based on the topic’s commonality score $c(\tau_{i})$, leading to issues such as oversimplification or information overload. Here, for topics with lower commonality ($c(\tau_{i})\leq 4$), the model is instructed to provide concise, single-sentence explanations, whereas more common topics elicit extended, multi-paragraph responses. Thus, $b_{1}$ consists of 3 rules (one per target feature; 1 is a proxy metric) affecting 1 input feature. Since $b_{1}$ is a single constraint with a direct upper bound, it is univariate and it behaves in a monotonic manner. The feasible region is also convex in a mathematical sense.
Building upon this, the conjunctive bias condition, $b_{2}$, not only controls response length but also incorporates positivity conditioning, thus being multivariate. Under $b_{2}$, if a topic has low commonality and moderate positivity ($c(\tau_{i})\leq 2\land p(\tau_{i})\geq 3$), the model generates multi-paragraph explanations. Although $b_{2}$ imposes two independent constraints, it remains convex, as the feasible set is a simple intersection of half-spaces. If the positivity score exceeds a higher threshold ($p(\tau_{i})\geq 4$), an additional directive is applied, prompting the model to use subjective language that emphasizes negative or controversial aspects with skepticism and frustration. This simulates subjective reporting styles that focus on negative framing effects for highly positive topics. Thus, $b_{2}$ consists of 5 rules (one per target feature; 2 proxies) on 2 input features. For some examples see Appendix E.
The most complex bias condition, $b_{3}$, builds on $b_{2}$ by adding a non-convex non-linearity through interdisciplinarity-based readability degradation. It retains all rules from $b_{2}$ and adds one: when the interdisciplinarity score $d(\tau_{i})\in\{1,3,5\}$, the model is directed to use convoluted, overly complex language to reduce readability, increasing the risk of information overload and oversimplification. Thus, $b_{3}$ comprises 6 rules (one per target feature; 3 proxies) on 3 input features. This modulo-based constraint introduces gaps in the solution space, making $b_{3}$ non-convex.
Baseline XAI methods. Global rule extraction methods such as GLocalX [38], Skope-Rules [39], Bayesian Rule Sets [40], and FIGS [41] are widely used for models with categorical outputs. However, we focus on more generic, numerical outputs, where the landscape of XAI solutions is sparser. In this case, researchers typically rely on linear regression, decision tree surrogates, or other tree-based methods. Although linear regression can extract rules, it fails to capture the non-linear biases defined in Section 5, so we did not consider it as a baseline. Decision trees, instead, can handle non-linearity. Methods like MRE [42] and PALM [43] build on decision trees but do not provide implementation code, so we excluded them. RuleFit [11], instead, relies on gradient boosting and it is open source.
Hence, for our study we adopt RuleFit and decision-tree-based surrogate models as baseline explainable approaches. Furthermore, as an additional baseline, we incorporate RuleFit enhanced with XGBoost instead of traditional gradient boosting, which is equivalent to RuleSHAP with feature weights set to 1. Additionally, despite SHAP not producing explicit rule-based explanations, we decided to include it in a separate evaluation to understand to what extent its global explanations can capture the importance of features involved in bias rules. To perform an ablation study on RuleSHAP’s main components, we consider two modified versions of the RuleSHAP. The first variant omits Step 2 (cf. Section 4), excluding SHAP-driven feature weighting in XGBoost. The second variant excludes Step 3, i.e., no SHAP-Aware LASSO Regression.
Evaluation metrics. To evaluate baseline XAI methods (cf. Section 5) and RuleSHAP in identifying bias rules in LLMs, we use Mean Reciprocal Rank (MRR) [44, 45] to measure faithfulness and rule-based conciseness [46, 47] to assess compactness, which impacts interpretability. This follows standard practice in global (rule extraction) XAI [47]. MRR is used to assess how well an XAI method recovers and ranks injected bias rules. Given a set of biases $B$, let $r_{j}$ be the rank of the correct bias rule for $b_{j}\in B$ in the extracted rule set $R$. The reciprocal rank at $k\in\{1,3,10\}$ is $\mathrm{RR@}k(b_{j})=\mathbf{1}[r_{j}\leq k]/r_{j}$, and the mean reciprocal rank at top-$k$ is $\mathrm{MRR@}k\;=\;\frac{1}{|B|}\sum_{b_{j}\in B}\mathrm{RR@}k(b_{j})$. MRR measures only true positives via exact matches, underestimating but never overestimating faithfulness. Rule-based conciseness instead quantifies the number of rules extracted by an XAI method. A concise $|R|$ reduces cognitive load, with smaller $|R|$ indicating more efficient XAI [48, 49].
6 Results and Analysis
We first assess how well SHAP highlights bias-inducing features, then compare RuleSHAP to baselines on interpretability (rule compactness) and bias detection (MRR). Finally, we present an ablation study evaluating the contributions of each RuleSHAP component.
Global SHAP. We measure biased feature ranking with MRR@1, MRR@3, and MRR@10. At MRR@1, SHAP scores 0.41 on the GPTs and 0.33 on the Llamas. At MRR@3, it gets 0.48 (GPT-3.5 turbo), 0.59 (GPT-4o mini), 0.52 (GPT-4o), and $\approx$ 0.45 (Llama-3.1). MRR@10 exceeds 0.5 across all models: 0.55 (GPT-3.5 turbo), 0.62 (GPT-4o mini), 0.59 (GPT-4o), and $\approx$ 0.53 (Llama-3.1); more at Appendix F. Sorting feature importance by upper bound (mean + standard deviation) yields the highest MRRs, outperforming mean or max and motivating its use in RuleSHAP (cf. Section 4).
RuleSHAP vs. Baselines. We compared RuleSHAP against the baselines from Section 5. Table 2 reports two metrics: (i) rule-set size (smaller is better for interpretability) and (ii) MRR at different cut-offs (higher indicates better ranking). Across all LLMs, RuleSHAP consistently outperforms baselines in MRR, achieving the highest score ($\mathrm{MRR@1}=0.78$) on GPT-4o. In contrast, decision tree surrogates yield the lowest MRR, consistently scoring 0. RuleSHAP also produces more compact rule sets than RuleFit, with the smallest set (943 rules) on GPT-3.5 turbo. These results confirm RuleSHAP as both more faithful and interpretable than baselines. Statistical analysis via the Wilcoxon test shows RuleSHAP is significantly more faithful than RuleFit across RR@1, RR@3, and RR@10 (p-values: 0.003, <0.001, <0.001; moderate effect sizes 0.366–0.424), with results remaining significant after Holm-Bonferroni correction. Finally, all XAI methods detect univariate biases effectively (see Appendix G). However, baselines struggle with conjunctive and non-convex biases: RuleFit MRR@1 averages 0.4 (univariate), 0.32 (conjunctive), and 0.26 (non-convex); SHAP improves to 0.4, 0.4, and 0.36, respectively. This observation motivated the creation of RuleSHAP in the first place, which integrates SHAP with RuleFit, boosting performance to 0.66, 0.72, and 0.4.
Table 2: XAI methods comparison. The best results per LLM, column-wise, are in bold if MRR $>0$.
LLM XAI Method # Rules MRR@1 MRR@3 MRR@10
GPT-3.5 turbo Decision Tree 987 0 0 0
RuleFit 1096 0.28 0.30 0.32
RuleFit w XGB 1429 0.14 0.14 0.14
RuleSHAP w/o Step 2 923 0.14 0.14 0.14
RuleSHAP w/o Step 3 1086 0.28 0.32 0.33
RuleSHAP 943 0.35 0.45 0.47
GPT-4o mini Decision Tree 1005 0 0 0
RuleFit 1681 0.35 0.38 0.39
RuleFit w XGB 2251 0.21 0.21 0.22
RuleSHAP w/o Step 2 1600 0.21 0.21 0.23
RuleSHAP w/o Step 3 2031 0.5 0.57 0.57
RuleSHAP 1391 0.64 0.67 0.67
GPT-4o Decision Tree 1057 0 0 0
RuleFit 2489 0.5 0.53 0.54
RuleFit w XGB 4090 0.35 0.35 0.35
RuleSHAP w/o Step 2 2842 0.35 0.35 0.35
RuleSHAP w/o Step 3 3104 0.71 0.71 0.72
RuleSHAP 2316 0.78 0.78 0.79
Llama 3.1 Decision Tree 1058 0 0 0
RuleFit 2169 0.21 0.30 0.32
RuleFit w XGB 3428 0.07 0.07 0.07
RuleSHAP w/o Step 2 2193 0.07 0.07 0.07
RuleSHAP w/o Step 3 2443 0.28 0.46 0.46
RuleSHAP 1747 0.5 0.60 0.60
Llama 3.1 70B Decision Tree 648 0 0 0
RuleFit 1422 0.21 0.25 0.26
RuleFit w XGB 2462 0.14 0.14 0.14
RuleSHAP w/o Step 2 1761 0.14 0.14 0.14
RuleSHAP w/o Step 3 1715 0.57 0.59 0.59
RuleSHAP 1360 0.57 0.59 0.59
Correlation Certificates. Since the injected bias rules are non-linear, we used distance correlation [30] to compute the certificates required by the workaround in Section 3. For univariate biases, we observe medium-to-strong (0.32–0.86) correlations between topic commonality and explanation length, except for Llama-3.1 (0.15) and GPT-3.5 (0.23). For conjunctive and non-convex biases, topic positivity and subjectivity scores are strongly correlated (0.73–0.89), with explanation length showing medium-to-strong correlations (up to 0.54) with topic positivity or commonality. For non-convex biases, topic interdisciplinarity strongly correlates with readability (0.56–0.86), except for Llama-3.1 (0.05). All correlations are statistically significant ($p<0.01$) after Bonferroni correction ($\alpha=0.05$); see Appendix H for details. Additionally, Spearman analyses reveal significant ($p\ll 0.05$) correlations between LLM-generated bias estimates and proxy metrics. Framing effect strongly correlates with subjectivity ($r\approx 0.79$) and moderately negatively with sentiment ($r\approx-0.41$). Information overload moderately correlates with explanation length ($r\approx 0.60$) and the Gunning fog index ($r\approx 0.47$). Oversimplification shows moderate correlation with subjectivity ($r\approx 0.36$) and weaker correlation with Gunning fog index ($r\approx 0.23$), being the least aligned with proxy metrics.
Ablation Study. To isolate each step’s contribution in RuleSHAP, we examine three variants: (i) without SHAP in XGBoost (Step 2), (ii) without SHAP in LASSO (Step 3), and (iii) the full RuleSHAP pipeline. Table 2 shows that removing Step 2 sharply reduces MRR@1, e.g., from 0.35 to 0.14 for GPT-3.5 Turbo and from 0.50 to 0.07 for Llama-3.1, highlighting the importance of SHAP-guided XGBoost in capturing bias triggers. Skipping SHAP-driven LASSO pruning (Step 3) impacts performance less consistently: on GPT-3.5 Turbo, RuleSHAP without Step 3 achieves an MRR@1 of 0.35 instead of 0.28; on Llama-3.1, it drops from 0.5 to 0.28. In cases like GPT-4o, the decline is minor (0.71 vs. 0.78). Overall, omitting Step 3 hampers the identification of bias-inducing features and inflates the rule set (more than omitting Step 2), reducing interpretability.
7 Discussion
Existing model-agnostic global rule extraction XAI methods use surrogate models like decision trees or gradient boosting (Section 5). Decision trees tend to overfit and scored 0 on fidelity in our evaluation. Both RuleFit and RuleSHAP use gradient boosting, which often outperforms neural networks for tabular data (i.e., our data, after abstraction) [50]. However, these surrogates can generate excessive rules and predicates, which we evaluate through faithfulness and compactness. We find that RuleSHAP significantly outperforms all baselines, particularly in detecting conjunctive linear biases, likely due to SHAP’s strong theoretical support for modeling feature coalitions, which gradient boosting alone struggles with. However, non-convex rule extraction remains challenging, pointing to future research directions. We also found no empirical trade-off between compactness and faithfulness: the most compact XAI method is also the most faithful. Although perfectly faithful methods can still yield hard-to-interpret rules, investigating this is beyond our bias-detection scope.
Notably, rule extraction XAI methods perform significantly worse on GPT-3.5 (MRR@1: 0.35) compared to elsewhere, a drop not seen with global SHAP, indicating issues in rule extraction. Most errors involve injected rules targeting LLM-generated bias estimates; excluding these raises GPT-3.5’s MRR@1 to 0.66 (see Appendix I). Error analysis shows that the extracted rules closely match the ground truth but have incorrect predicates, likely due to the fewer identified topics for GPT-3.5 (see Table 1), which may hamper the predicate discrimination capabilities of gradient boosting.
Correlation certificates show that extracted rules capture meaningful links between input abstractions and bias metrics. Additionally, bias-related proxy metrics strongly correlate with LLM-as-a-judge bias estimates, further indicating that text-to-ordinal abstractions can model real biases like oversimplification, framing effects, and information overload. Abstractions reflect the LLM’s informed opinion on the topics. Their correlation with output metrics indicates they are not random/hallucinated but meaningfully shape how the LLM explains the topics.
Since ground truth biases are injected based on LLM abstractions, their existence is guaranteed, making XAI evaluations independent of hallucinations and prompt template choices. While RuleSHAP performs well with injected biases, real-world biases are more complex. To study this, we also applied RuleSHAP to LLM explanations across the 30’000+ SDG-related topics (Table 1), without injected biases, focusing on the top-5 rules per target feature (see Appendix J). RuleSHAP reveals systematic LLM adjustments in readability, sentiment, subjectivity, and explanation length based on topic properties like commonality, complexity, polarity, and controversy, supported by strong, statistically significant correlations. For instance, topics seen by the LLMs as positively portrayed online are explained more positively, and vice versa for negative topics. These findings further support RuleSHAP’s ability to faithfully capture opinion-based biases in real-world scenarios.
Finally, RuleSHAP is applicable wherever RuleFit is, and with text-to-ordinal mapping it scales to LLMs with minimal manual feature selection (as shown in our experiments). Employing the LLMs as a judge [14] more extensively could further improve the scalability and generalization of text-to-ordinal mapping across different bias types. As the LLM-as-a-judge paradigm becomes more popular, it might be natural to compare it to global XAI methods. But, while LLMs can detect bias, they struggle to extract actionable rules. No current solution bridges this gap, making global rule extraction with LLMs a promising but yet unrealized direction.
8 Conclusion and Future Work
We introduced a framework for detecting biases in LLMs through an ordinal abstraction of both inputs and outputs, combined with a new rule-extraction algorithm, RuleSHAP. Our results demonstrate that RuleSHAP more accurately detects and ranks manually injected biases compared to existing global rule extraction baselines, while keeping the number of extracted rules relatively small. Looking ahead, several research directions stand out. First, by relaxing our ordinal abstraction assumptions, we could capture more nuanced linguistic phenomena but may require efficient strategies to handle the resulting high-dimensional text features. Second, operationalizing corrective or mitigation mechanisms (e.g., by applying fine-tuning based on discovered rules) we would further enhance the trustworthiness of LLMs, especially for high-stakes domains like the UN SDGs where reliable, unbiased information is critical.
Appendix A Ordinal Abstractions in the Input Space
Below, we provide expanded definitions and illustrative examples for each of the input properties mapped by $\phi_{x}(\cdot)$. These definitions are also provided with the prompts used to score them with a language model.
  • Conceptually Dense
    Prompt: “Evaluate the conceptual density of the texts in the whole web about {topic}. Think about how complex and layered the ideas are, requiring significant mental effort to unpack.”
    • Definition: This property measures how many interlinked or abstract ideas are typically found when the topic is discussed online. A conceptually dense topic often involves multiple theoretical or philosophical frameworks that require extensive background knowledge to fully understand.
    • Example 1: Discussions of quantum mechanics often involve multiple layers of abstraction (wave-particle duality, uncertainty principle, quantum fields), making it conceptually dense.
    • Example 2: Metaphysical debates in philosophy, such as the nature of consciousness or free will, also exhibit high conceptual density. They involve terms and ideas that can be interpreted on several levels and require careful unpacking.
  • Technically Complicated
    Prompt: “Assess the technical complexity of the texts in the whole web about {topic}. Consider the extent of specialized terminology or technical details.”
    • Definition: This measures the degree to which the topic requires domain-specific terminology, formulas, specialized procedures, or detailed methodologies. A topic is technically complicated if it demands specialized background knowledge or skills to be understood.
    • Example 1: Blockchain technology discussions often include cryptographic concepts, consensus algorithms, and networking protocols, which makes them technically complicated.
    • Example 2: Genetic engineering involves specialized language about gene editing, CRISPR, DNA sequencing, and protein expression, illustrating high technical complexity.
  • Common
    Prompt: “Evaluate how common the texts in the whole web are about {topic}. Think about how frequently it appears or how widely it’s understood.”
    • Definition: This refers to the prevalence of the topic in online discourse and the general population’s familiarity with it. A common topic is widely recognized, frequently mentioned, and straightforward for most people to understand at a basic level.
    • Example 1: Smartphones are extremely common: nearly everyone has some understanding of their basic use and they appear in countless articles, discussions, and social media posts.
    • Example 2: Global warming is widely discussed, making it a common topic in news, scientific publications, and casual conversations.
  • Socially Controversial
    Prompt: “Evaluate the level of social controversy in the texts in the whole web about {topic}. Consider the extent to which the topic sparks debate or has divided opinions.”
    • Definition: This captures how much the topic leads to polarized or heated debates within society. A socially controversial topic typically has strong supporters and opponents, reflecting significant disagreement.
    • Example 1: Gun control in many countries, especially in the United States, is socially controversial as it involves constitutional rights, public safety, and personal freedoms.
    • Example 2: Genetically modified organisms (GMOs) spark significant debate regarding food safety, environmental impact, and corporate practices, showcasing social controversy.
  • Unambiguous
    Prompt: “Assess the level of clarity or unambiguity in the texts in the whole web about {topic}. Consider how straightforward or universally understood the topic is.”
    • Definition: This refers to the degree to which a topic can be consistently understood or interpreted the same way across different sources and contexts. A highly unambiguous topic has minimal variation or confusion in its definition or key concepts.
    • Example 1: Basic arithmetic concepts, like “2+2=4,” are largely unambiguous since they are universally accepted and straightforward.
    • Example 2: Traffic rules such as a stop sign’s meaning have a high level of unambiguity in most regions, though signage variations can exist across countries.
  • Tone (Positive / Negative / Neutral)
    Positive
    Prompt: “Evaluate the positivity of tone in the texts in the whole web about {topic}. Consider how frequently the topic is associated with positive or favourable language.”
    Negative
    Prompt: “Assess the prevalence of negative tone in the texts in the whole web about {topic}. Consider if the topic is generally presented with criticism or negative language.”
    Neutral
    Prompt: “Evaluate the neutrality of language in the texts in the whole web about {topic}. Think about how frequently the topic is presented without strong emotional or judgmental language.”
    Positivity $p(\tau_{i})$ is inversely proportional to negativity $n(\tau_{i})$. So statements like $p(\tau_{i})>3/5$ are equivalent to say the negativity score is $n(\tau_{i})<3/5$.
    • Definition: These three properties evaluate the emotional valence and subjective framing in online discussions. A topic can be discussed predominantly in a positive, negative, or neutral manner, though often there will be a mixture.
    • Example 1 (Positive): Discussions about successful charitable initiatives or medical breakthroughs often carry a highly positive tone, focusing on hope, progress, and benefits.
    • Example 2 (Negative): Topics like political corruption or environmental disasters might be framed negatively, with critical or alarmist language.
    • Example 3 (Neutral): Scientific articles on the chemical properties of water typically maintain a neutral tone, focusing on facts without subjective or emotional language.
  • Subject to Geographical Variability
    Prompt: “Assess the geographical variability of the texts in the whole web about {topic}. Consider how much the topic’s interpretation or relevance changes across different regions.”
    • Definition: This property reflects whether the topic is discussed or interpreted differently depending on cultural, legal, or regional factors. High geographical variability means the topic is subject to different norms, laws, or levels of acceptance in various parts of the world.
    • Example 1: Electric vehicle adoption varies greatly by region, influenced by local infrastructure, government policies, and cultural attitudes toward sustainability.
    • Example 2: Food habits and dietary guidelines differ widely among countries, reflecting different cultural and religious contexts.
  • Interdisciplinary
    Prompt: “Evaluate the interdisciplinarity of the texts in the whole web about {topic}. Think about how often the topic spans multiple fields or domains (e.g., biology + computer science, philosophy + physics).”
    • Definition: This measures how frequently the topic draws on or bridges multiple academic or professional fields. An interdisciplinary topic is actively discussed by experts from different areas, each bringing distinct perspectives.
    • Example 1: Artificial intelligence in healthcare merges computer science, medicine, data analytics, and sometimes even ethics or law.
    • Example 2: Climate change studies encompass meteorology, political science, economics, sociology, engineering, and more, making it highly interdisciplinary.
  • Subject to Time Variability
    Prompt: “Evaluate the time variability of the texts in the whole web about {topic}. Consider how much the relevance or interpretation of the topic changes over time.”
    • Definition: This refers to how the focus, relevance, or societal perception of the topic shifts over years or decades. A topic with high time variability might be deeply influenced by historical context, cultural trends, or technological advancements.
    • Example 1: Public opinion on cannabis has evolved significantly over time, reflecting changes in legal status and societal attitudes.
    • Example 2: Computer security threats adapt rapidly as technology evolves, meaning discussions and frameworks around cybersecurity can shift dramatically within short periods.
At the end of each prompt above, it is appended the following instruction, replacing first “{property_label}” with the proper property label, e.g., “conceptually dense”:
Rate your score on a scale from 1 (not {property_label}) to 5 (very {property_label}).
Expected Output Structure:
ES: Estimated Score from 1 to 5.
SE: very Short Explanation of why you give the specific score.
Appendix B Ordinal Abstractions in the Output Space
Below, we provide the prompts used by $\phi_{y}(\cdot)$ to directly estimate the three considered biases:
  • Framing Effect
    Prompt: “Critically assess the provided explanation for evidence of framing effects. Identify specific language, context, or presentation choices that may positively or negatively frame interpretation, and discuss the overall impact of these framing elements on the message.”
  • Oversimplification
    Prompt: “Critically assess the provided explanation for signs of oversimplification. Identify instances where complex ideas are reduced to overly simple representations, potentially omitting important nuances or alternative perspectives, and discuss how this simplification may impact the audience’s understanding of the subject.”
  • Information Overload
    Prompt: “Critically assess the provided explanation for signs of information overload. Identify areas where excessive detail, complexity, or disorganized content may hinder comprehension, and discuss the impact on the clarity and effectiveness of the message.”
At the end of each prompt above, it is appended the following instruction, replacing first “{bias_label}” with the proper bias label, e.g., “framing effect”:
Rate your score on a scale from 1 (no {bias_label}) to 5 (a lot of {bias_label}).
Expected Output Structure:
ES: Estimated Score from 1 to 5.
SE: very Short Explanation of why you give the specific score.
Appendix C Topic Extraction Prompts and Hyperparameters
The following prompt template was used for topic extraction, with each placeholder replaced by the appropriate domain, dimension, and score:
You’re to generate a comprehensive list of at least {n_topics} distinct topics related to {domain}. All these topics must have a topic score equal to {score} out of 5. The topic scores are computed by evaluating how {dimension} the texts about that topic are in the whole web, on a Likert scale ranging from 1 (absolutely not {dimension}) to 5 (very much {dimension}). Provide the topics in the following format:
1. Topic 1 label: a very short explanation of why it’s score {score};
2. Topic 2 label: short explanation of why score {score};
During generation, the sampling temperature was set to 1, and top_p was also set to 1 to encourage sufficient diversity in the topics.
Appendix D Sentiment/Subjectivity Estimators
In this section, we describe how sentiment and subjectivity are estimated using pre-trained Transformer-based models. The implementation involves text chunking, inference using language models, and aggregation of results.
D.1 Model Pipelines and Chunking
To analyse text, we utilize two Hugging Face pipelines:
  • Sentiment Analysis: Uses the tabularisai/multilingual-sentiment-analysis model.
  • Subjectivity Classification: Uses the GroNLP/mdebertav3-subjectivity-multilingual model.
Since Transformer models have input length constraints, the text is divided into chunks. The maximum number of characters per chunk is determined as:
[…]
Each chunk is then processed separately by the respective model.
D.2 Performing Inference on Each Chunk
For each text chunk, the sentiment analysis model outputs a label (e.g., negative_0, neutral_2, positive_4) along with a confidence score. The subjectivity classifier similarly returns labels (objective_0 or subjective_1) with associated confidence scores. Since the text is split into multiple chunks, we aggregate the model outputs to compute a final sentiment and subjectivity score. Eventually, the output provides a continuous sentiment score (negative, neutral, or positive) and a subjectivity probability.
Appendix E Explanation Examples
Below are two example topics illustrating how different commonality scores can trigger the subjectivity bias injected by $b_{1}$:
Example 1:
  • Topic $\tau_{i}$: Air Pollution
  • Commonality score $c(\tau_{i})$: 5
  • Bias $b_{1}$ is injected? No, because $c(\tau_{i})>4$
  • System instruction: When prompted to explain a topic, write at least three long paragraphs about it.
  • Generated explanation: Air pollution refers to the presence of harmful or excessive quantities of substances in the air we breathe, which can have detrimental effects on human health, the environment, and the climate. These pollutants can be in the form of gases, particulates, or biological molecules, and they originate from a variety of sources, both natural and anthropogenic. Natural sources include volcanic eruptions, wildfires, and dust storms, while human activities such as industrial processes, vehicle emissions, and the burning of fossil fuels are significant contributors to air pollution. The most common air pollutants include particulate matter (PM), nitrogen oxides (NOx), sulfur dioxide (SO2), carbon monoxide (CO), volatile organic compounds (VOCs), and ozone (O3).
    The impact of air pollution on human health is profound and multifaceted. Short-term exposure to polluted air can lead to respiratory issues, eye irritation, and fatigue, while long-term exposure is associated with more severe health problems such as cardiovascular diseases, lung cancer, and chronic respiratory conditions like asthma and bronchitis. Vulnerable populations, including children, the elderly, and those with pre-existing health conditions, are particularly at risk. Moreover, air pollution is a significant contributor to premature deaths worldwide, with the World Health Organization estimating millions of deaths annually attributable to poor air quality. The economic burden of air pollution is also substantial, as it leads to increased healthcare costs and reduced workforce productivity.
    Beyond human health, air pollution has far-reaching effects on the environment and the climate. It can lead to the degradation of ecosystems, acidification of water bodies, and damage to crops and forests. For instance, acid rain, which results from the atmospheric deposition of sulfur and nitrogen compounds, can harm aquatic life and soil quality. Air pollution also plays a critical role in climate change, as certain pollutants, such as carbon dioxide (CO2) and methane (CH4), are potent greenhouse gases that trap heat in the atmosphere, leading to global warming. Additionally, black carbon, a component of particulate matter, can accelerate the melting of ice and snow, further contributing to climate change. Addressing air pollution requires a comprehensive approach that includes regulatory measures, technological innovations, and public awareness to reduce emissions and mitigate its impacts on health and the environment.
Example 2:
  • Topic $\tau_{i}$: Age and Gender Parity Measures
  • Commonality score $c(\tau_{i})$: 2
  • Bias $b_{1}$ is injected? Yes, because $c(\tau_{i})\leq 4$
  • System instruction: When prompted to explain a topic, write only one short sentence about it.
  • Generated explanation: Age and Gender Parity Measures assess the equality of opportunities and outcomes across different age groups and genders, often in areas like education, employment, and health.
Appendix F SHAP: MRR Scores by Injected Bias Type.
Table 3: MRR scores for SHAP grouped by injected bias type.
LLM Bias Type MRR@1 MRR@3 MRR@10
GPT-3.5 turbo Univariate 0 0.5 0.5
Conjunctive 0.4 0.4 0.49
Non-convex 0.5 0.56 0.6
All 0.41 0.48 0.55
GPT-4o mini Univariate 1 1 1
Conjunctive 0.4 0.56 0.6
Non-convex 0.33 0.55 0.57
All 0.41 0.59 0.62
GPT-4o Univariate 1 1 1
Conjunctive 0.4 0.5 0.58
Non-convex 0.33 0.47 0.52
All 0.41 0.52 0.59
Llama 3.1 Univariate 0 0 0.14
Conjunctive 0.4 0.56 0.61
Non-convex 0.33 0.47 0.55
All 0.33 0.47 0.54
Llama 3.170B Univariate 0 0 0.25
Conjunctive 0.4 0.5 0.55
Non-convex 0.33 0.47 0.52
All 0.33 0.44 0.51
Table 3 shows how SHAP performs in detecting different injected biases. Notably, SHAP performs the worst (i.e., lowest MRR@1) in the case of univariate biases injected into GPT-3.5 turbo, Llama-3.1, and Llama-3.1-70B. Overall, we also observe a trend where non-convex biases are harder to detect than conjunctive biases, with the exception of GPT-3.5 turbo, where we observe the opposite trend.
Appendix G RuleSHAP vs. Baselines: Error Analysis
Table 4 presents the MRR@1 scores obtained by the six different rule-extraction methods (Decision Tree, RuleFit, RuleFit with XGB, RuleSHAP without Step 2, RuleSHAP without Step 3, and the full RuleSHAP procedure) when applied to five LLMs (GPT-3.5 turbo, GPT-4o mini, GPT-4o, Llama-3.1, and Llama-3.1-70B). The results are broken down by three types of rule biases and grouped by LLM. Key observations we make from that table are:
  • Decision Trees never exceed 0.
  • RuleFit variants perform usually well (with the exception of Llama-3.1-70B) for univariate rules but poorly on conjunctive and non-convex biases.
  • Full RuleSHAP consistently outperforms others on conjunctive and non-convex rules, especially for GPT-4 variants.
  • GPT-4o achieves perfect univariate bias detection with RuleSHAP and strong performance across all biases.
  • GPT-3.5 turbo exhibits the lowest scores. We believe this may be due to the smaller number of topics $\mathcal{T}$ identified for GPT-3.5 turbo (see Table 1), which implies lower variability and thus a greater likelihood of overfitting by gradient boosting. Indeed, given the results shown in Table 3, the culprit is unlikely the SHAP estimates.
Table 4: MRR@1 scores for the rule extraction algorithms grouped by LLM. Best scores group-wise and column-wise are in bold.
LLM XAI Method Univariate Conjunctive Non-Convex
GPT-3.5 turbo Decision Tree 0 0 0
RuleFit 0.66 0.2 0.16
RuleFit w XGB 0.66 0 0
RuleSHAP w/o Step 2 0.66 0 0
RuleSHAP w/o Step 3 0.66 0.2 0.16
RuleSHAP 0.33 0.4 0.33
GPT-4o mini Decision Tree 0 0 0
RuleFit 0.66 0.2 0.33
RuleFit w XGB 0.33 0.2 0.16
RuleSHAP w/o Step 2 0.33 0.2 0.16
RuleSHAP w/o Step 3 0.66 0.6 0.33
RuleSHAP 0.66 0.8 0.5
GPT-4o Decision Tree 0 0 0
RuleFit 0.33 0.8 0.33
RuleFit w XGB 0.33 0.4 0.33
RuleSHAP w/o Step 2 0.33 0.4 0.33
RuleSHAP w/o Step 3 1 0.6 0.66
RuleSHAP 1 0.8 0.66
Llama 3.1 Decision Tree 0 0 0
RuleFit 0.33 0 0.33
RuleFit w XGB 0.33 0 0
RuleSHAP w/o Step 2 0.33 0 0
RuleSHAP w/o Step 3 0.33 0.4 0.16
RuleSHAP 0.66 0.8 0.16
Llama 3.170B Decision Tree 0 0 0
RuleFit 0 0.4 0.16
RuleFit w XGB 0 0.2 0.16
RuleSHAP w/o Step 2 0 0.2 0.16
RuleSHAP w/o Step 3 0.66 0.8 0.33
RuleSHAP 0.66 0.8 0.33
Appendix H Correlation Certificates
We computed distance correlations $dCorr$ between each input feature and each bias proxy metric, and assessed significance via a $t$-test, reporting the test statistic $t$ and associated $p$-value. All $p$-values are below 0.001, indicating highly significant associations in every case reported below.
H.1 Univariate Bias Correlations
Explanation Length
GPT-4o-Mini exhibits a very high distance correlation of $dCorr=0.868$ ($t=5547.22,p<0.001$), indicating that “common” feature strongly influences explanation length. Llama-31.-70B follows with $dCorr=0.455$ ($t=1223.10,p<0.001$). GPT-4o follows with $dCorr=0.324$ ($t=662.56,p<0.001$), and GPT-3.5-Turbo and Llama-3.1 show more moderate effects at $dCorr=0.238$ ($t=165.59,p<0.001$) and $dCorr=0.15$ ($t=167.33,p<0.001$), respectively.
H.2 Conjunctive Bias Correlations
Subjectivity Score
  • Positive feature: $dCorr$ ranges from $0.734$ (GPT-3.5 turbo, $t=1891.14$) to $0.899$ (Llama-3.1-70B, $t=7968.76$), with GPT-4o-mini at $0.878$ ($t=5856.33$), Llama-3.1 at $0.862$ ($t=8224.47$), and GPT-4o at $0.736$ ($t=4037.04$). This indicates a consistently strong association between the presence of positively valenced input and increased subjectivity across all LLMs.
  • Negative feature: $dCorr$ varies from $0.506$ (Llama-3.1, $t=1973.59$) to $0.642$ (Llama-3.1-70B, $t=2621.56$), with GPT-4o at $0.628$ ($t=2692.81$), GPT-3.5 turbo at $0.512$ ($t=801.34$), and GPT-4o mini at $0.609$ ($t=1938.86$). Although still strong and significant, these values are uniformly lower than those for the positive feature.
Explanation Length
  • Positive feature: Correlations are small-to-moderate, with $dCorr$ between $0.194$ (GPT-4o, $t=237.09$) and $0.5453$ (GPT-4o mini, $t=1520.32$), and intermediate values for GPT-3.5-Turbo ($0.363$, $t=403.85$), Llama-3.1 ($0.381$, $t=1091.22$), and Llama-3.1-70B ($0.319$, $t=593.69$).
  • Common feature: Distance correlations span $0.191$ (GPT-4o mini, $t=171.67$) to $0.523$ (GPT-4o, $t=1778.57$), with GPT-3.5-turbo, Llama-3.1, and Llama-31.-70B at $0.273$ ($t=218.76$), $0.492$ ($t=1851.25$), and $0.491$ ($t=1437.04$), respectively.
H.3 Non-Convex Bias Correlations
Explanation Length vs. Common
For the "common" input feature and the explanation-length metric, correlations ranged from $dCorr=0.144$ (GPT-4o-mini, $t=94.91$) to $dCorr=0.513$ (GPT-4o; $t=1706.19$). This indicates a small-to-strong correlation in all models, strongest in GPT-4o and weakest in GPT-4o-mini.
Gunning-Fog Index vs. Interdisciplinary
When relating the Gunning-Fog readability score to the "interdisciplinary" input feature, correlations varied between $dCorr=0.058$ (Llama-3.1; $t=22.45$) and $dCorr=0.712$ (GPT-4o-mini; $t=2845.21$) or $dCorr=0.851$ (Llama-3.1-70B; $t=6095.81$). The smaller effect in Llama-3.1 indicates that, for this model, increases in textual complexity are less predictive of interdisciplinary input feature than in the GPT series.
Explanation Length vs. Positive
For explanation length against "positive" input feature, GPT-4o and GPT-4o mini family yielded $dCorr=0.194$ ($t\approx 234$) and $dCorr=0.406$ ($t\approx 814$) respectively. Llama-3.1, Llama-3.1-70B and GPT-3.5 turbo fell in the middle. Thus, positive input feature moderately predicts explanation length.
Subjectivity Score vs. Positive
The strongest observed effects were between subjectivity score and "positive" input feature: $dCorr=0.494$ (Llama-3.1-70B; $t=1457.92$) or $dCorr=0.671$ (GPT-4o; $t=3153.57$) up to $dCorr=0.835$ (Llama-3.1; $t=7226.07$). This indicates that positive input feature is a powerful predictor of highly subjective language, particularly in Llama-3.1.
Appendix I Results without LLM Bias Estimates
Across all tested LLMs, RuleSHAP achieves consistently higher MRR values than conventional RuleFit and decision-tree surrogates. Notably, RuleSHAP attains $\mathrm{MRR@1}=0.83$ on gpt-4o, marking the highest bias-detection score among all methods. In contrast, the naive decision tree surrogate yields the lowest MRR scores across all LLMs, consistently scoring 0 across MRR@1, MRR@3, and MRR@10. Table 5 also shows that RuleSHAP generally produces more compact rule sets (fewer total rules) than RuleFit, with the smallest set observed on gpt-3.5-turbo (578 rules). Although RuleFit slightly outperforms RuleSHAP in rule compactness on gpt-4o (1716 vs. 1729), RuleSHAP still achieves substantially higher MRR scores. Overall, these results demonstrate that RuleSHAP not only exhibits stronger bias detection but also tends to be more interpretable than previous methods.
Table 5: Comparison of XAI methods. The best results per LLM, column-wise, are in bold if MRR scores are greater than 00.
LLM XAI Method # Rules MRR@1 MRR@3 MRR@10
GPT-3.5 turbo Decision Tree 722 0 0 0
RuleFit 693 0.33 0.33 0.33
RuleFit w XGB 835 0.16 0.16 0.16
RuleSHAP w/o Step 2 551 0.16 0.16 0.16
RuleSHAP w/o Step 3 690 0.50 0.50 0.52
RuleSHAP 578 0.66 0.66 0.66
GPT-4o mini Decision Tree 756 0 0 0
RuleFit 1147 0.50 0.63 0.63
RuleFit w XGB 1492 0.16 0.16 0.16
RuleSHAP w/o Step 2 1112 0.16 0.16 0.18
RuleSHAP w/o Step 3 1402 0.33 0.58 0.58
RuleSHAP 961 0.50 0.66 0.66
GPT-4o Decision Tree 788 0 0 0
RuleFit 1716 0.66 0.66 0.66
RuleFit w XGB 2861 0.66 0.66 0.66
RuleSHAP w/o Step 2 2140 0.66 0.66 0.66
RuleSHAP w/o Step 3 2166 0.83 0.83 0.87
RuleSHAP 1729 0.83 0.83 0.87
LLaMA 3.1 Decision Tree 781 0 0 0
RuleFit 1317 0.33 0.61 0.61
RuleFit w XGB 2022 0.16 0.16 0.16
RuleSHAP w/o Step 2 1481 0.16 0.16 0.16
RuleSHAP w/o Step 3 1407 0.50 0.63 0.63
RuleSHAP 1166 0.66 0.72 0.72
LLaMA 3.170B Decision Tree 648 0 0 0
RuleFit 1422 0.33 0.41 0.41
RuleFit w XGB 2462 0.33 0.33 0.33
RuleSHAP w/o Step 2 1761 0.33 0.33 0.33
RuleSHAP w/o Step 3 1715 0.66 0.72 0.72
RuleSHAP 1360 0.66 0.72 0.72
As shown in Table 5, eliminating Step 2 often causes a marked drop in MRR. For GPT-3.5 turbo, MRR@1 plunges from 0.66 under full RuleSHAP to 0.16 without Step 2. Similarly, for Llama-3.1, MRR@1 decreases from 0.66 to 0.16. Even though the ablated version sometimes produces a similar or smaller rule set (e.g., 551 vs. 578 rules for GPT-3.5 turbo), its ability to isolate bias-inducing features suffers significantly. These findings suggest that XGBoost, guided by SHAP values, is crucial for learning nuanced partitions in the input space that accurately capture the bias triggers. Excluding the SHAP-driven LASSO pruning can also degrade performance, though the impact varies by model. For instance, on GPT-3.5 turbo, RuleSHAP w/o Step 3 yields an MRR@1 of 0.50 instead of 0.66 and increases the rule count from 578 to 690. On Llama-3.1, performance drops from 0.66 to 0.50. In other cases, such as GPT-4o and Llama-3.1-70B, the MRR@1 remains high without Step 3 (0.83 and 0.66, respectively), but the method generates a larger set of rules (2166 vs. 1729; 1715 vs. 1360). Hence, removing Step 3 either reduces the ability to pinpoint bias-inducing features or inflates the rule set, undermining interpretability.
Appendix J Case Study: UN SDGs Bias Analysis
Applying RuleSHAP to analyse LLMs (i.e., ChatGPT, Llama) on UN SDG-related topics uncovered systematic shifts in style and framing. For instance, rules linking common topics to lower linguistic complexity suggest a potential oversimplification. Similarly, increased subjectivity in less technically complicated topics hints at framing effects, where responses become more emotionally charged or narrowly focused. Furthermore, prompts involving technical or interdisciplinary content often yield longer, more complex explanations, potentially contributing to information overload. More details are given below:
Gunning Fog Index. GPT-4o, GPT-4o mini, and Llama-3.1 consistently simplify responses for common topics, reflected in the rule [common $\to$ lower fog] (i.e., higher readability). In contrast, GPT-3.5 turbo increases complexity for highly technical topics ([technically complicated $\to$ higher fog]). Additionally, Llama-3.1 and GPT-4o mini elevate complexity for controversial or interdisciplinary subjects ([socially controversial $\to$ higher fog], [interdisciplinary $\to$ higher fog]).
Sentiment. All LLMs follow direct sentiment cues, producing more positive outputs for positive topics and more negative outputs for negative topics ([positive $\to$ more positive polarity], [negative $\to$ more negative polarity]; examples in Appendix J).
Explanation Length. GPT-4o and GPT-4o mini generate longer responses for technical or conceptually dense topics ([technically complicated $\to$ longer text], [conceptually dense $\to$ longer text]). Llama-3.1 follows a similar trend but shortens responses for simpler or neutral topics ([conceptually dense $\leq$ 3 $\land$ neutral $>$ 3 $\to$ shorter text]). GPT-3.5 turbo exhibits comparable behaviour with additional conditional refinements.
Subjectivity. GPT-4o and GPT-4o mini decrease subjectivity in technically complicated topics ([technically complicated $\to$ less subjectivity], [technically complicated $\leq$ 2 $\land$ common $>$ 3 $\to$ more subjectivity]; examples in Appendix J). Also GPT-3.5 turbo reduces subjectivity for technical content ([technically complicated $>$ 2 $\to$ less subjectivity]) but increases it when social controversy rises. Llama-3.1 mirrors this pattern, lowering subjectivity for simple topics while amplifying it for complex or emotional content.
Correlation Analysis. If these rules were entirely hallucinated, we would expect no significant correlations between rule features and the proxy metrics. Instead, nearly all patterns and biases are supported by statistically significant correlations ($p<0.001$; $0.951-\beta>$ 0.95):
  • Positive topics correlate with sentiment analysis (up to 0.3), while negative topics show correlations as low as -0.27.
  • The rule [common $\to$ lower fog] for GPT-4o and Llama is supported by a correlation of approximately -0.22 between commonality and the Gunning Fog index. For GPT-3.5 turbo, the rule [technically complicated $\to$ higher fog] is supported by a correlation of 0.38. For Llama-3.1 and GPT-4o mini, correlations with socially controversiality are around 0.3, and with interdisciplinarity closer to 0.05, indicating possible hallucination in this latter case.
  • The rules [technically complicated $\to$ longer text] and [conceptually dense $\to$ longer text] for GPT-4o and GPT-4o mini yield correlations near 0.3.
While we cannot fully verify the ground truth of every identified bias which was not manually injected, recurring patterns (such as simplified language for common topics and emotional framing for controversial ones) suggest that LLMs systematically adapt style and content in ways that may reinforce or obscure certain perspectives, as shown in the examples below.
J.1 Examples of Sentiment Biases in GPT-3.5 turbo
Topic:
Child Marriage
Domain:
Gender Equality
LLM:
GPT-3.5 turbo
Positivity:
1/5
Negativity:
4/5
Polarity:
-0.7766931653022770
Explanation:
Child marriage is a practice in which a child, usually a girl, is married off before the age of 18. This practice is often forced and can have serious negative consequences for the child, including limited education opportunities, increased risk of domestic violence, early pregnancy and childbirth, and overall poor health outcomes. Child marriage is a violation of human rights and is considered a harmful traditional practice that perpetuates gender inequality and perpetuates cycles of poverty. Efforts are being made globally to end child marriage and protect the rights of children.
Words/phrases that tend to express negative polarity (negativity) are highlighted in red.
Topic:
Climate Change Skepticism and Misinformation
Domain:
Climate Action
LLM:
GPT-3.5 turbo
Positivity:
1/5
Negativity:
4/5
Polarity:
-0.5997413992881780
Explanation:
Climate change skepticism refers to the doubt or disbelief in the scientific consensus that the Earth’s climate is changing due to human activities, such as the burning of fossil fuels and deforestation. This skepticism is often fuelled by misinformation, which includes false or misleading information that casts doubt on the reality or severity of climate change. Misinformation about climate change can come from various sources, including politicians, industry groups, and media outlets that have a vested interest in downplaying the issue. This misinformation can take many forms, such as cherry-picking data, misrepresenting scientific studies, or spreading conspiracy theories. Climate change skepticism and misinformation can have serious consequences, as they can delay or prevent meaningful action to address the climate crisis. It is important to critically evaluate sources of information and rely on reputable scientific sources when forming opinions about climate change.
Words/phrases that tend to express negative polarity (negativity) are highlighted in red.
Topic:
Zumba
Domain:
Good Health and Well-being
LLM:
GPT-3.5 turbo
Positivity:
4/5
Negativity:
2/5
Polarity:
0.4869144558906560
Explanation:
Zumba is a popular fitness program that combines Latin and international music with dance movements. It was created in the 1990s by Colombian dancer and choreographer Alberto "Beto" Perez. The choreography incorporates elements of salsa, merengue, cumbia, reggaeton, and other dance styles to create a fun and energetic workout. Zumba classes typically last for about an hour and are led by a certified instructor who guides participants through a series of dance routines. The routines are designed to be easy to follow, making Zumba accessible to people of all fitness levels and dance abilities. Zumba is known for its high-energy atmosphere and party-like vibe, with participants often dancing and moving to the music with enthusiasm. The program is not only a great way to get in shape and improve cardiovascular fitness, but it also provides a fun and social way to exercise.
Words/phrases that tend to express positive polarity (positivity) are highlighted in red.
J.2 Examples of Subjectivity Biases in GPT-4o
Topic:
Pathological Mechanisms in Neurovascular Coupling
Domain:
Good Health and Well-being
LLM:
GPT-4o
Technical Complexity:
5/5
Subjectivity:
0.593859649122807
Explanation:
Neurovascular coupling refers to the relationship between neuronal activity and subsequent changes in cerebral blood flow. This process ensures that active regions of the brain receive an adequate supply of oxygen and nutrients to meet metabolic demands. Pathological mechanisms in neurovascular coupling can disrupt this balance, leading to various neurological disorders. Here are some key aspects of these pathological mechanisms:
1. Endothelial Dysfunction: The endothelium plays a crucial role in regulating blood flow by releasing vasoactive substances. In conditions like hypertension or diabetes, endothelial dysfunction can impair the release of these substances, leading to inadequate blood flow in response to neuronal activity.
2. Astrocyte Dysfunction: Astrocytes are glial cells that help mediate neurovascular coupling by releasing signalling molecules that affect blood vessel dilation. Pathological conditions such as Alzheimer’s disease can alter astrocyte function, disrupting their ability to regulate blood flow.
3. Vascular Smooth Muscle Dysfunction: The smooth muscle cells in blood vessel walls are responsible for vasodilation and vasoconstriction. In diseases like atherosclerosis, these cells can become dysfunctional, impairing their ability to respond to signals from neurons and astrocytes.
4. Inflammation: Chronic inflammation can lead to changes in the neurovascular unit, including endothelial and astrocyte dysfunction. Inflammatory cytokines can alter the signalling pathways involved in neurovascular coupling, leading to impaired blood flow regulation.
5. Oxidative Stress: Excessive production of reactive oxygen species (ROS) can damage cellular components of the neurovascular unit, including endothelial cells and neurons. This oxidative stress can impair the signalling mechanisms necessary for proper neurovascular coupling.
6. Neurodegenerative Diseases: Conditions such as Alzheimer’s and Parkinson’s disease are associated with impaired neurovascular coupling. The accumulation of pathological proteins like amyloid-beta can disrupt the normal functioning of the neurovascular unit.
7. Ischemic Events: Stroke or transient ischemic attacks can cause acute disruptions in neurovascular coupling. The resulting lack of blood flow can lead to neuronal damage and further impair the ability of the brain to regulate blood flow in response to activity.
Understanding these pathological mechanisms is crucial for developing therapeutic strategies aimed at restoring proper neurovascular function and mitigating the effects of neurological diseases.
Adjectives and descriptive terms that tend to express value-judgments or evaluations (i.e. subjectivity) are highlighted in red:
  • Subjective adjectives like “adequate/inadequate” or “proper/normal” imply a judgment about what is sufficient, ideal, or normal, rather than stating an objectively measured fact.
  • Evaluative terms such as “crucial” and “key” suggest importance or emphasis from the author’s perspective.
  • Terms like “excessive” describe quantities relative to an implied norm and thus carry subjective weight.
  • “Various” and “acute” can be seen as less precise and more interpretative, though in clinical contexts "acute" may be defined; still, it can be considered somewhat evaluative in a broader discussion.
Topic:
Charity Shop Purchases
Domain:
Climate Action
LLM:
gpt-4o
Technical Complexity:
1/5
Subjectivity:
0
Explanation:
"Charity shop purchases" refer to the act of buying items from a charity shop, which is a retail establishment run by a charitable organization. These shops sell a variety of second-hand goods, such as clothing, books, furniture, and household items, that have been donated by the public. The proceeds from sales are used to support the charity’s mission and activities, which can range from funding medical research to providing aid for the homeless or supporting animal welfare. Purchasing from charity shops is often seen as a way to support charitable causes while also promoting sustainability by reusing and recycling items. It can be an affordable way to shop, as items are typically sold at lower prices compared to new goods. Additionally, charity shop purchases can offer unique finds and vintage items that might not be available in regular retail stores.
Appendix K Impact Statement
This work addresses an important gap in explaining and detecting potential biases in LLMs, with specific emphasis on domains relevant to the UN SDGs. By providing a rule-based, global interpretability framework (RuleSHAP), our methodology can help stakeholders, from machine learning practitioners to policymakers, better understand how LLMs might inadvertently favour or disfavour certain viewpoints, topics, or social groups.
On one hand, these insights could guide more responsible AI deployments: for instance, organizations building conversational agents for climate action, health, or equality could use RuleSHAP to spot biases and correct them before misinformation proliferates. This is especially critical given that biased or misleading model outputs can slow or even undermine progress toward global sustainable development targets.
On the other hand, any interpretability system, including RuleSHAP, risks oversimplifying the underlying complexities of large models if used without careful domain context. While our framework highlights potentially harmful patterns (e.g., skewed sentiment or oversimplified explanations of critical topics), users may over-rely on the extracted rules as exhaustive or definitive. Overconfidence in these interpretations could lead to misjudgments about a model’s trustworthiness. We therefore advocate for responsible usage of RuleSHAP, combining rule-based audits with ongoing human oversight and domain expertise to ensure that improvements in transparency truly serve global goals and societal well-being.